OpenStudy (anonymous):

2√6 + 3√26 how will i do this simplified form

OpenStudy (anonymous):

sorry for askign the question i got the answer

OpenStudy (anonymous):

Is not that in simplified form already???

OpenStudy (anonymous):

it can't be further simplified

OpenStudy (anonymous):

It can be simplified but that will look weird.... @dirtymind ..

OpenStudy (anonymous):

\[2 \sqrt{6} + 3 \sqrt{26} \implies 2 \sqrt{3 \times 2} + 3 \sqrt{13 \times 2} \implies \color{green}{\sqrt{2}[2 \sqrt{3} + 3 \sqrt{13}]}\]

OpenStudy (anonymous):

I have colored it so that it does not look weird...